During 2019, the standard deviation of ACT scores of the students at UA was 40. A sample of 16 students had a standard deviation of 45. Testing to see if there has been a significant change in the standard deviation of ACT scores at UA.

a. Develop null and alternative hypotheses.

b. What is the value of test statistic?

c. At 95% confidence, what are the critical values (C.V)

d. What is the conclusion?
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: σ = 40
Alternative Hypothesis, Ha: σ ≠ 40

Test statistic,
Χ^2 = (n-1)*s^2/σ^2
Χ^2 = (16 - 1)*45^2/40^2
Χ^2 = 18.984

This is two tailed test, for α = 0.05 and df = 15
Critical value of Χ^2 are 6.262 and 27.488

Do not reject null hypothesis
There is no significant change in the standard deviation